OpenStudy (thadds2003):

Sylvia has a total of 16 trading cards: 8 superhero cards, 4 princess cards, 2 villain cards, and 2 wildlife cards. What is the ratio of superhero cards to the total number of trading cards in simplest form?

OpenStudy (thadds2003):

@Michele_Laino

OpenStudy (anonymous):

1:2

OpenStudy (anonymous):

since it is 8 to 16

OpenStudy (anonymous):

medal?

OpenStudy (michele_laino):

hint: total number of cards is \(16+8+4+2+2=...\)

OpenStudy (thadds2003):

= 32.

OpenStudy (anonymous):

it is 16 trading cards

OpenStudy (anonymous):

see what I mean

OpenStudy (michele_laino):

sorry I have made a typo: total number of cards is: \(8+4+2+2=...?\)

OpenStudy (thadds2003):

16

OpenStudy (anonymous):

there are 32 cards in total but 16 of them are trading cards

OpenStudy (anonymous):

see what I mean

OpenStudy (anonymous):

medal?

OpenStudy (michele_laino):

so, the requested ration is: \(8/16=...?\), since the number superhero cards is 8

OpenStudy (thadds2003):

so thats the answer?

OpenStudy (anonymous):

why do you have to be 13 for QH?

OpenStudy (anonymous):

why not 12 since a ton of the population is young

OpenStudy (michele_laino):

you have to divide both numerator and denominator by 8

OpenStudy (thadds2003):

1 : 2.

OpenStudy (anonymous):

divide them both by 8 which is the farthest you can go

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

medal?

OpenStudy (michele_laino):

so: \(8:8=...?\) \(16:8=...?\)

OpenStudy (anonymous):

cause 1 is not a prime but can't be divided anymore and 2 is a prime

OpenStudy (thadds2003):

1 : 2

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

will this site be around in a year 1/2

OpenStudy (anonymous):

a year and a half?

OpenStudy (anonymous):

thx for medal

OpenStudy (thadds2003):

yes.

OpenStudy (anonymous):

phew cause I am going to be 13 then

OpenStudy (anonymous):

making me ready for QH

OpenStudy (anonymous):

meaning I would be Qualified to be a Qualified Helper get it?

OpenStudy (anonymous):

any other questions

OpenStudy (anonymous):

if you do link them in another question

OpenStudy (anonymous):

but I do have a 14 year old sis